Qadir, Raj, and Sabrina go together to a local video store. Qadir rents two movies and three games for a total cost of $24.30. Raj rents three movies and one game for a total cost of $18.25. How much money does Sabrina need to rent one movie and two games?
Answer by elima4(15)

Q = 2m + 3g = 24.30
R = 3m + g = 18.25
Now we subtract the equations to find out one of the variables, put first we need to multiply one of the equations with a number so when they are subtracted it eliminates one of the variables. We will multiply 3 to R equation;
(3)3m+g = 18.25
9m+3g = 54.75
Now we can subtract;
2m+3g=24.3
-(9m+3g=54.75)
---------------
-7m=-30.45
m=4.35
So we know the movies are $4.35
Lets plug that into the either equation to find g;
2(4.35)+3g=24.30
8.7+3g=24.30
3g=15.6
g=5.2
Games = $5.20
S = 4.35+2(5.2)
S = 4.35 + 10.4
S = 14.75
:)